A model of a helicopter rotor has four blades, each 3.40 m

long from the central shaft to the blade tip. The model is rotated in a wind tunnel at 550 rev/min (a) What is the linear speed of the blade tip, in (b) What is the radial acceleration of the blade tip expressed as a multiple of the acceleration of gravity, g?


1
Expert's answer
2022-03-04T08:32:31-0500

Answer

Size or radius of blade

R=3.40m

Angular speed

"w=57.60rad\/sec"

a) linear speed

"v=Rw=3.40*57.60\\=195.84m/s"

b) radial acceleration of the blade tip

"a=\frac{v^2}{R}\\=\frac{195.84*195.84}{3.40}\\=11280.38m/s^2"